OpenStudy (anonymous):

Solve 5x2 - 7x + 2 = 0 by completing the square. What is the constant added on to form the perfect square trinomial? 49/4 49/25 49/100

OpenStudy (anonymous):

\[5x^{2} - 7x +2 = 0\]\[5(x^{2} - \frac{7}{5}x) + 2 = 0\]\[5(x^{2} - \frac{7}{5}x + \frac{49}{100} - \frac{49}{100}) + 2 = 0\]\[(x-\frac{7}{10})^{2} - \frac{49}{20} + 2 = 0\]\[(x-\frac{7}{10})^{2} - \frac{9}{20} = 0\]

OpenStudy (jhannybean):

Really good job.

OpenStudy (anonymous):

what would i do after your last step?

OpenStudy (anonymous):

@Jhannybean thank you! :)

OpenStudy (jhannybean):

That is the final step. otherwise youd be solving for x and that involves a multitude of other steps.

OpenStudy (anonymous):

yep!

OpenStudy (dumbcow):

no its says to solve for "x" using completing the square -you want to isolate the squared term, then take sqre root

OpenStudy (jhannybean):

oh you DO want to solve haha ah......

OpenStudy (anonymous):

oh yes, sorry i read that wrong! thanks for correcting it @dumbcow

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

\[(x-\frac{7}{10})^{2} = \frac{9}{20}\]\[x = \sqrt{\frac{9}{20}}+\frac{7}{10}\]

OpenStudy (jhannybean):

\[(x-\frac{7}{10})^{2} - \frac{9}{20} = 0\]\[(x-\frac{7}{10})^2 = \frac{9}{20}\]sqare root both sides\[\sqrt{(x-\frac{7}{10})^2}=\sqrt{\frac{9}{20}} \]\[x-\frac{7}{10} = \pm \frac{3}{\sqrt{20}}\]\[x= \frac{7}{10} \pm \frac{3}{20}\]

OpenStudy (jhannybean):

\[ x= \frac{7}{10} \pm \frac{3}{2\sqrt{5}}\]

OpenStudy (jhannybean):

i forgot the square root on my second to last post.

OpenStudy (anonymous):

which answer would i pick then ?

OpenStudy (dumbcow):

here are some general formulas to use: given a quadratic equation \[ax^{2} +bx +c = 0\] completing the square yields \[a(x+\frac{b}{2a})^{2} +(c-\frac{b^{2}}{4a}) = 0\] solving gives the quadratic formula \[x = \frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (jhannybean):

the "constant" added to coplete the square would be the value you'd obtain by finding the value "c". You have to fit your equaion into the quadratic form \[\large ax^2 +bx+c=0\] and to find your value of c, you evaluate your value of "b" :- \[\large c : \large \left(\frac{-7/5}{2}\right)^2 = \frac{49}{100}\]

OpenStudy (anonymous):

i thought that i was supposed multiply the 5 to the 2 (c term) and then subtract that number to both sides. then after that dont i take half of the negative seven and then put it into x+3.5

OpenStudy (anonymous):

thank you @Jhannybean

OpenStudy (jhannybean):

Yeah...what you're talking about can also be solved like this.you don't necessarily HAVE to complete the square. I'll show you a different method.

OpenStudy (jhannybean):

\[\large 5x^2 -7x+2=0\]\[\large \color{green}5x^2-7x+\color{green}2=0\]\[\large x^2 - 7x + 10 =0\]\[\large (x-5)(x-2)=0\]\[\large (x-\frac55)(x-\frac25)=0\] Simplify whatever you can. \[\large (x-1)(5x-2)=0 \] Solve for x's. \\[\large x-1 =0 \]\[\large 5x-2 =0 \]\[\large x = 1 \ ,\ \frac25\]

OpenStudy (jhannybean):

Gotta go! -Jhannybean takes leave-

OpenStudy (anonymous):

isnt that finding the factors though?

OpenStudy (jhannybean):

Mmhmm. That is. it's just another method. :D

OpenStudy (anonymous):

ah ok thank you :D

OpenStudy (jhannybean):

np :D